Splitz & Joinz manual

Daniel Pfund September 18, 1995

version 1.1 \begin{matrix}c \\ ◯ \end{matrix}1995 Daniel Pfund

Contents

1 Introduction 1 2 Distribution 2 3 Uses 2 4 Installation 2 5 Memory requirements 3 6 Running the program 3 7 How the program works 3

7.1 Splitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 7.2 Joining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

8 Limitations and bugs 4 9 Legal stuff 4 10 Author 4

Summary This file contains all the necessary information to use and install

Splitz & Joinz .

1 Introduction

Splitz & Joinz is a program designed to split and join text files. The program is written in OPL and will only work on a Psion3a.

2 Distribution

Splitz & Joinz is now freeware. This means you can distribute and use it freely. The program can be distributed by any electronic mean (anonymous FTP sites, BBSes, packet radio …) provided the file “SPLITZ.ZIP” stays intact. Software distributors can freely distribute this program on their shareware disks, provided that:

1. You don’t charge more than $5 for the service.

2. The file “SPLITZ.ZIP” stays intact (see section 4 “Installation” for the file’s contents).

3. You don’t make any money out of it.

If you would like to put this program on a CD-ROM, please contact me first (see section 10 “Author” for my address).

3 Uses

I originally made it to avoid the 40k limit of the inbuilt applications like Word and OPL. It will process text files and split them “intelligently” after a LineFeed (LF) at about the size you specify. I used it for big text files I received while I had to use my Psion3a for packet radio (a radio transmission mode for radio amateurs) while my Amiga’s serial port was defectuous. You might also want to use it to split/join large text files if you are writing big articles on your Psion…

4 Installation

You should unzip the file SPLITZ.ZIP and get the following files:

  1. SPLITZ.TXT This file

2. SPLITZ.DVI This file in L^{A}TEXDVI format

3. SPLITZ.OPA The program

All you have to do is copy the file SPLITZ.OPA to the \APP directory with whatever program you use to copy/move files from/to your Psion. You will then have to install the program with the System menu (Apps/Install) or by pressing Psion-i. Since the program doesn’t use any specific files (ie: it is a type $1000 application), you can also rename the file “SPLITZ.OPA” to “SPLITZ.OPO” and put that file into your RunOPL directory (usually this is “\OPO”. This will avoid using one more icon on your System screen and allow you to group all little utilities under one icon.

5 Memory requirements

Splitz & Joinz will need only about 3k of your disk space. In addition, Splitz & Joinz will need about 28k of free memory to be able to run. Also, don’t forget that you need space for the output of the file you want to split/join.

6 Running the program

You will first get a simple menu with 3 possibilities. The descriptions are rather self explaining, so make your choice! Once you have chosen the “Start” option, you will be prompted with the main dialog where you are requested to enter useful information. The first line asks for the source filename and it’s location (Disk A, Internal, Disk B, eventually REM::). You can use Tab or Control-Tab to get the usual Psion file requesters. The third line asks for the destination directory. The default is “LOC::M\”, although it will only show you “\”. Once again, you can specify which drive the destination directory should be on. The fifth line asks you for the approximate size of the parts. This is only used if you split a file of course. Splitz & Joinz will automatically detect if you want to split or join a file by looking at the source filename. If the source file has an extension of the kind “P0 ?”, it will attempt to join the files previously splitted. Otherwise it will take the source file and attempt to split it. Note that the minimal split value is 8k. When joining, Splitz & Joinz will automatically ask you for the new filename extension. When you enter it, you do not need to write a “.” in the dialog, but only the last three letters. The last line of the main dialog asks you if you want to delete the source file after use. If you answer yes, Splitz & Joinz will delete the source file, but only if there were no errors during the process of splitting/joining. Once you have entered everything, type “Enter” to start splitting/joining. If everything goes well, you will get a dialog saying “No errors” and eventually another line telling you which files have been deleted. If on the other hand, an error appeared during the process of splitting/joining, you will be informed by a requester and the program will end without deleting any files.

7 How the program works

7.1 Splitting

The program checks if it can read 8192 bytes (8k) of the source file without going over the split value (hence the minimal split value of 8k ; for the sake of speed). If it can, it reads 8k and writes them out to the destination file ; otherwise it reads one byte and checks if this character is a LF. If it is, it writes it to the destination file and closes it. It checks if there are any bytes left in the source file and if yes, creates a new part. The process then starts over.

7.2 Joining

Joining is much simpler in the way that it takes all the files with the same name and extensions “P0 ?”, starting at 1 and joins them together. Note that in the main dialog, you do not necessarily need to choose part1, as it will recognize any part.

8 Limitations and bugs

• The minimal split value is set to 8k, otherwise the program would be somewhat slow.

• You cannot split a program into more than 9 parts, since the extensions are of the kind “P0 ?”.

• No bugs have been found during testing, although I have never tried to run the program on a binary file.

9 Legal stuff

Splitz & Joinz is provided “as is”. This means, it does what it does, and I cannot be held responsable if it crashes your system, or causes any other dammages. If however you do have a problem with it, please report it to me and I will try to help you out to fix it.

10 Author

Snail mail: Daniel Pfund 19, ch. de Tirelonge CH-1213 Onex Switzerland

You can also reach me via E-Mail (preferred !): “Pfund@uni2a.unige.ch” or if this address doesn’t work any more (in a couple of years) you can still write to me at: “hb9vbc@pccr01.cern.ch”. (Note the zero, not an o !)

[End of Splitz & Joinz documentation]